An Object Moves along the X-Axis. Its X-Coordinate is given as a Fuction of time as x = 7t-3t2
Where x is in Metres and t is in Seconds. Its Average Speed over the Interval t=0 to t=4 is ?
 
(a) 5m/s
(b) -5m/s
©  -169/24m/s
(d) 169/24m/s
 
Give Answer with Detailed Explaination if Possible. Thanks.

Grade:11

3 Answers

Yash Puthran
9 years ago
Anybody
 
siddharth gupta
9 years ago
POSTION OF THE OBJECT AT DIFFERENT INSTANTS OF TIME ARE:
at t=1,x=4 and distance travelled=4m
at t=2,x=2 and distance travelled=6m  (4+2)
at t=3,x= – 6 and distance travelled=14 m (6+8)
at t=4,x= – 20 and distance travelled=28 m (14+14)
average speed=(total distance)/(total time)=28/4=7m/s=169/24m/s(as it is speed it will always be positive).
WITH REGARDS
SIDDHARTH GUPTA
Sher Mohammad IIT Delhi
9 years ago
POSTION OF THE OBJECT AT DIFFERENT INSTANTS OF TIME ARE:
at t=1,x=4 and distance travelled=4m
at t=2,x=2 and distance travelled=6m  (4+2)
at t=3,x= – 6 and distance travelled=14 m (6+8)
at t=4,x= – 20 and distance travelled=28 m (14+14)
average speed=(total distance)/(total time)=28/4=7m/s=169/24m/s(as it is speed it will always be positive